OpenStudy (anonymous):

Describe how to transform the quantity of the fifth root of x to the seventh power, to the third powerinto an expression with a rational exponent. Make sure you respond with complete sentences.

OpenStudy (ybarrap):

$$ \large{ \left(\left (\sqrt[5]{x}\right )^7\right)^3\\ =\left (\left (x^{1/5}\right )^7\right))^3\\ =\left(x^{7/5}\right)^3\\ =x^{\frac{7*3}{5}}\\ =x^{\frac{21}{5}}\\ } $$ Expand the 5th root of x into an exponent with a fraction, 1/5 In the exponent of x, multiply 7 times this value In the exponent of x, Multiply the result times 3

OpenStudy (anonymous):

wouldnt the 7+3=10 instead of 11?

OpenStudy (nikato):

It's actually 7x3 which =21

OpenStudy (anonymous):

so its x 21 over 5?

OpenStudy (nikato):

Yes. x^ 21/5

OpenStudy (ybarrap):

@JTsoccer Good question. That is a common mistake. You are thinking about something like x^7 * x^3 = x^ (7+3) = x^(10) However, (x^7)^3 is not the same as x^7 * x^3. Because, (x^7)^3 = x^7 * x^7 * x^7 = x^(7+7+7) = x^(21).

Directrix (directrix):

@pizzaqueen5